OpenStudy (blackops2luvr):

-7x + y = -12 7x + 5y = -18

OpenStudy (babyslapmafro):

7x + (-7x) = 5y + y = -12 + (-18) =

OpenStudy (babyslapmafro):

Add each equation with respect to the variables

OpenStudy (blackops2luvr):

-30

OpenStudy (blackops2luvr):

wat

OpenStudy (blackops2luvr):

wait

OpenStudy (babyslapmafro):

yes -12 + (-18) = -30

OpenStudy (blackops2luvr):

0 5 -30

OpenStudy (babyslapmafro):